3. Evaluate determinants of (SH cos no sin nel 1-sin no cos nel 114 | 2 15 2 51 o 8 8 -21
Answer #1

3.6 I 14 2 5 2 5 0 8 8 – 2 12 ol 8 = 14 (o-by) - 2 (-4-40) +5 (16-0) = - 14 x64 + 2x44 +5816 -896 # 88 + 80 2 – 728 Am. cosno